OpenStudy (anonymous):

Simplify. 2 ^3√108 1) 16 ^3√18 2) 8 ^3√6 3) 5 ^√4 4) 6 ^3√4

OpenStudy (anonymous):

So which option?

myininaya (myininaya):

\[2\sqrt[3]{108}=2\sqrt[3]{3^3*4}=2*3\sqrt[3]{4}=6\sqrt[3]{4}\]

myininaya (myininaya):

marc i told you this will happen no one is going to understand your notation

myininaya (myininaya):

rethink your notation lol

myininaya (myininaya):

2*(108)^(1/3) is fine

OpenStudy (anonymous):

Haha i know im trying to but idk how to make the notation when i ask a question

myininaya (myininaya):

\[2\sqrt[3]{108}\]=2*(108)^(1/3) ok?

OpenStudy (anonymous):

Yes...How do i make the √ go over the whole number? what do i press

myininaya (myininaya):

\[\sqrt[?]{?}\] to do this just put sqrt[ ]{ }

OpenStudy (anonymous):

sqrt[4]{x}

myininaya (myininaya):

try typing sqrt[3]{108} into equation editor below

myininaya (myininaya):

do you see equation below?

OpenStudy (anonymous):

I can do that...but how do i do that and ask a question? theres no equation button

myininaya (myininaya):

i guess you can always type the question after making the post that you need help